John's age is three years more than twice Bill's age. Three years ago John was four times as old as Bill was then. How old is each now?
Answer by checkley77(12844)

NOW:
JOHN=X+3
BILL=X
3 YEARS AGO:
JOHN=(X+3-3)
BILL=(X-3)4
(X+3-3)=4(X-3)
X=4X-12
X-4X=-12
-3X=-12
X=-12/-3
X=4 BILL'S AGE NOW.
4+3=7 JOHN'S AGE NOW.
PROOF:
(4+3-3)=(4-3)4
4=1*4
4=4